# CADmium: Fine-Tuning Code Language Models for Text-Driven Sequential CAD Design

### v1

@@ Appendix C Prompt Details
To evaluate the quality of our CADmium annotations against Text2CAD expert annotations, we employed an LLM-as-a-judge setup using Gemma-3 12B [23], as described in Section 3.1 and with results presented in Figure 2b. In this evaluation, 10,000 paired annotations (one from CADmium, one from Text2CAD expert for the same CAD model) were assessed. To mitigate positional bias, the CADmium and Text2CAD annotations were randomly assigned to the labels “1.” and “2.” within the user message presented to Gemma-3 12B. For optimizing the classification process and minimizing output variability, the model was explicitly instructed to respond with only a single digit (0, 1 or 2), allowing the judgment to be captured efficiently. The specific prompts used to instruct Gemma-3 12B for each of the four evaluation criteria are detailed as follows:
@@ Appendix D Qualitative Analysis of New Metrics
Appendix E Licenses for External Assets
@@
The DeepCAD dataset [1] is available under the MIT License, as indicated in its official repository. The ABC dataset [12], which is a source for DeepCAD, is also distributed under the MIT License. The Fusion 360 Reconstruction dataset [10], originating from the Autodesk Fusion 360 Gallery, is referenced in academic literature as being available for research purposes under the Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) license.
The language models leveraged include: Qwen2.5-Coder [7], open-source versions of which are provided under the Apache License 2.0. The GPT-4.1 model [6] is a proprietary offering from OpenAI, and its usage is governed by OpenAI’s prevailing Terms of Use, Usage Policies, and Sharing & Publication Policy. Gemma models [23], developed by Google, are released under a custom Gemma license.
Key software utilized includes Blender [22], which is distributed under the GNU General Public License (GPL) version 3 or later. The Python library trimesh [28], used for geometric computations, is available under the MIT License.
@@
Appendix F Licenses for New Assets

### v2

@@ Appendix C Prompt Details
To evaluate the quality of our CADmium annotations against Text2CAD expert annotations, we employed an LLM-as-a-judge setup using Gemma-3 12B [Team et al., 2025], as described in Section 3.1 and with results presented in Figure 2b. In this evaluation, 10,000 paired annotations (one from CADmium, one from Text2CAD expert for the same CAD model) were assessed. To mitigate positional bias, the CADmium and Text2CAD annotations were randomly assigned to the labels “1.” and “2.” within the user message presented to Gemma-3 12B. For optimizing the classification process and minimizing output variability, the model was explicitly instructed to respond with only a single digit (0, 1 or 2), allowing the judgment to be captured efficiently. The specific prompts used to instruct Gemma-3 12B for each of the four evaluation criteria are detailed as follows:
@@ Appendix D Qualitative Analysis of New Metrics
Appendix E Comparitive Qualitative Analysis
Appendix F Licenses for External Assets
@@
The DeepCAD dataset [Wu et al., 2021] is available under the MIT License, as indicated in its official repository. The ABC dataset [Koch et al., 2019], which is a source for DeepCAD, is also distributed under the MIT License. The Fusion 360 Reconstruction dataset [Willis et al., 2021], originating from the Autodesk Fusion 360 Gallery, is referenced in academic literature as being available for research purposes under the Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) license.
The language models leveraged include: Qwen2.5-Coder [Yang et al., 2024], open-source versions of which are provided under the Apache License 2.0. The GPT-4.1 model [OpenAI, 2025] is a proprietary offering from OpenAI, and its usage is governed by OpenAI’s prevailing Terms of Use, Usage Policies, and Sharing & Publication Policy. Gemma models [Team et al., 2025], developed by Google, are released under a custom Gemma license.
Key software utilized includes Blender [Blender Online Community, 2025], which is distributed under the GNU General Public License (GPL) version 3 or later. The Python library trimesh [Dawson-Haggerty et al., ], used for geometric computations, is available under the MIT License.
@@
Appendix G Licenses for New Assets